I can not figure out this math problem and need help. The question says, "Bill wnats 30 pounds of a coffee blend that he can sell for $7.00 per pound. If he has coffee that sells for $6.00 a pound and another brand that sells for $7.50 per pound, how many pounds of each brand should he use to make his blend?"
Answer by ptaylor(2198)

Let x=number of pounds of the $6.00 brand
Then 30-x=number of pounds of $7.50 brand

Now we know that the amount of money generated by the $6.00 brand ($6.00)(x) plus the amount of money generated by the $7.50 brand (30-x)($7.50) must equal the amount of money generated by the final blend (30)($7.00) so our equation to solve is:

6.00(x)+7.50(30-x)=(30)(7.00) simplifying we have:
6x-7.5x+225=210 collecting like terms we get:
-1.5x=-225+210=-15
x=10 lb for the $6.00 brand
30-x=30-10=20 lb for the $7.50 brand
ck
6.00(10)+7.50(30-10)=30(7.00)
60+150=210
210=210
Hope this helps. Happy holidays.--------ptaylor